Circular No. 7409

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

PROBABLE NOVA IN AQUILA Kesao Takamizawa, Saku-machi, Nagano, Japan, reports his discov-

ery of an apparent nova (mag 10.0) on unfiltered 5-min T-Max 400 film exposures taken on Apr. 28.669 and 28.673 UT with twin 0.10-m f/4 patrol cameras, giving the star’s position as α = 18^{h}52^{m}02^{s}.0, δ = −1^{o}28′14′′ (equinox 2000.0). Nothing appears on Takamizawa’s earlier patrol films from 1994 May 12 to 2000 Apr. 11 (limiting mag 13.5–15.0). T. Kato, Kyoto University, reports that Minoru Yamamoto, Aichi, independently discovered this object at mag 9.8 on a T-Max 400 exposure taken on Apr. 28.694. Visual magnitude estimates forwarded by Kato: Apr. 29.654, 9.8 (S. Takahashi, Shiga); 29.708, 10.0 (H. Hirosawa, Aichi); 29.731, 9.3 (H. Maehara, Saitama). It is uncertain if this star is identical with the suspected 1917 nova (peak m_{pg} ∼ 11) and eclipsing binary (IBVS 4232, 4338) CI Aql, which has position end figures 03^{s}.66, 38′′.7 (via Reinmuth 1925, A.N. 225,

385) and appears identical with an R ∼ 15.7 red star that is present in the USNO-A2.0 star catalogue (position end figures 03^{s}.56, 38′′.9).

M. Uemura and T. Kato, Kyoto University, report that they obtained low-dispersion spectra (range 430–730 nm) of the new star on Apr. 29.6 UT with the Okayama Astrophysical Observatory 1.88-m telescope. Hα emission, whose FWHM is ∼ 2300 km/s, is prominent and Hβ is also detected. The spectrum is similar to those of fast novae a few weeks after maximum. The Balmer lines show no P-Cyg profiles. These spectral features indicate that the outbursting object is a fast nova at a slightly evolved stage.

SUPERNOVA 2000ca IN ESO 383-32 J. Maza, University of Chile; and M. Hamuy, University of Arizona,

report the discovery by R. Antezana, on a 30-min unfiltered T-Max 400 film taken by L. González on Apr. 28.18 UT with the University of Chile’s Maksutov telescope, of an apparent supernova (m_{pv} ∼ 16.0) located at α = 13^{h}35^{m}22^{s}.98, δ = −34^{o}09′37′′.0 (equinox 2000.0), which is 0′′.6 east and 4′′.7 north of the nucleus of ESO 383-32 (z = 0.024). SN 2000ca was confirmed by Hamuy and A. Zuniga from BVRI images taken with the Cerro Tololo 0.9-m telescope on Apr. 29.27, and by M. Roth from JHK images taken with the Las Campanas 1.02-m Swope telescope on Apr. 29.3. From these images, estimated magnitudes for SN 2000ca are B = 15.91 ± 0.05, V = 15.99 ± 0.05, which suggests that the supernova is in an early stage. A T-Max film taken on Apr. 7.263 showed nothing at the position of the new object (limiting mag ∼ 19.0).

2000 April 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green